OpenStudy (anonymous):

What am I doing wrong? "The canister shown measures 7" in diameter and has a height of 7.5". Find the capacity (volume). Do not round your answer." The formula for Volume in a cylinder is the base times the height. I ended up with 164.85, but the answer was supposed to be 288.4875

OpenStudy (anonymous):

area of a cylinder= (base)*(height) which is =(area of circle)*(height) so area of cylinder=\[(\pi r^2) (height)\] ****remember radius is half of diameter*** so r=3.5 finally, \[\pi * 3.5^2*7.5=288\]